# Exploring Graph-Transformer Out-of-Distribution Generalization Abilities

### v1

@@ Appendix A Experimental details
Dataset Dist. shift MPNN GT #Layers #Params Batch size #Max epochs LR Weight decay
GOOD-SST2 Length vGIN 4 4.3M 32 200 1e-3 0
GOOD-SST2 Length MHA 5 3.8M 32 200 1e-3 0
GOOD-SST2 Length GIN MHA 10 3.8M 32 200 1e-3 0
GOOD-CMNIST Color vGIN 3 2.7M 128 500 1e-3 1e-4
GOOD-CMNIST Color MHA 4 2.9M 128 500 1e-3 1e-4
GOOD-CMNIST Color GIN MHA 7 2.7M 128 500 1e-3 1e-4
GOOD-Motif Size vGIN 4 3.8M 32 200 1e-3 0
GOOD-Motif Size MHA 5 3.6M 32 200 1e-3 0
GOOD-Motif Size GIN MHA 10 3.6M 32 200 1e-3 0
GOOD-Motif Basis vGIN 4 3.8M 32 200 1e-3 0
GOOD-Motif Basis MHA 5 3.6M 32 200 1e-3 0
GOOD-Motif Basis GIN MHA 10 3.6M 32 200 1e-3 0
@@
We perform graph-level prediction experiments on three datasets, across four covariate shift splits, using seven DG algorithms and three backbone methods, as detailed in Table 3. Each group of three consecutive rows corresponds to the evaluated backbones: vGIN for the MPNN, MHA for the GT, and GPS as the hybrid. Our implementation builds upon PyTorch Geometric [51] and the GOOD benchmark framework [8]. To ensure fair comparison, the number of parameters is kept roughly similar across backbones for each task, with GPS often having fewer parameters.
For experiments on SST2 and Motif, we use the Adam optimizer with no weight decay. In the CMNIST experiments, we adopt smaller models (with the number of layers specified in Table 3) and use the AdamW optimizer with a weight decay of 0.0001, which led to better performance. Across all experiments, we apply a dropout rate of 0.5, ReLU activation, and mean global pooling. The hidden layer size is set to 300. Batch size, maximum number of training epochs, and learning rates are listed in Table 3. As part of a preliminary analysis, we found that five-dimensional RW embeddings offer greater robustness than the ten-dimensional version.
@@ Appendix B In-distribution experimental results discussion
Figure 6 provides a comprehensive visualization of ID test accuracy results across the four benchmark datasets. This analysis complements our OOD findings by validating that high OOD performance does not come at the expense of staying competitive—or even superior—under standard ID conditions. These results also appear in Table 1. On CMNIST-color, GPS improves the vGIN results significantly across all DG algorithms, while MHA notably decreases ID accuracy. On Motif task, all backbones overall perform comparably. On SST2-length vGIN performs slightly better than MHA and GPS. These findings are crucial for practical applications where both ID and OOD samples are encountered, as they show that the superior OOD generalization capabilities of GPS hybrid do not come at the expense of standard ID performance.

### v2

@@ Appendix A Experimental details
Dataset / Shift MPNN GT #Layers #Params BSZ #Epochs LR WD Scheduler
SST2-Length vGIN 3 1.7M 32 200 1e-3 0 cosine
SST2-Length MHA 5 3.8M 32 200 1e-3 0 step 150
SST2-Length GIN MHA 3 2.9M 32 200 1e-3 0 cosine
CMNIST-Color vGIN 7 2.7M 128 500 1e-3 1e-4 step 300
CMNIST-Color MHA 4 2.9M 128 500 1e-3 1e-4 step 300
CMNIST-Color GIN MHA 3 2.7M 128 500 1e-3 1e-4 step 300
Motif-Size GIN 4 1.2M 32 200 1e-3 0 none
Motif-Size MHA 5 3.6M 32 200 1e-3 0 step 150
Motif-Size GIN MHA 4 3.6M 32 200 1e-3 0 step 150
Motif-Basis GIN 4 1.2M 32 200 1e-3 0 none
Motif-Basis MHA 5 3.6M 32 200 1e-3 0 step 150
Motif-Basis GIN MHA 4 3.6M 32 200 1e-3 0 step 150
HIV-Size vGIN 3 1.5M 32 200 1e-3 0 cosine
HIV-Size MHA 2 1.5M 32 200 1e-3 0 cosine
HIV-Size GIN MHA 3 1.8M 32 200 1e-3 0 cosine
HIV-Scaffold vGIN 3 1.5M 32 200 1e-3 1e-4 step 150
HIV-Scaffold MHA 2 1.5M 32 200 1e-3 1e-4 step 150
HIV-Scaffold GIN MHA 2 1.8M 32 200 1e-3 1e-4 step 150
@@
We perform graph-level prediction experiments on three datasets, across six covariate shift splits, using seven DG algorithms and three backbone methods, as detailed in Table 3. Each group of three consecutive rows corresponds to the evaluated backbones: vGIN/GIN for the MPNN, MHA for the GT, and GPS as the hybrid. Our implementation builds upon PyTorch Geometric [53] and the GOOD benchmark framework [8]. To ensure a fair comparison, we initially matched the number of parameters across backbones for each task. However, when varying the number of layers, we observed that in some cases smaller models achieved higher accuracy, leading us to select model sizes accordingly.
For experiments on CMNIST and HIV-Scaffold, we use the AdamW optimizer with 1e-4 weight decay, which led to better performance. Across all experiments, we apply a dropout rate of 0.5, ReLU activation, and mean global pooling. The hidden layer size is set to 300. Batch size (BSZ), maximum number of training epochs, and learning rates (LR) are listed in Table 3. Two schedulers were used: cosine-annealing with 10-epoch warmup and cosine decay to 10% LR, and single-step, which decays LR by 0.1 at specified milestone. As part of a preliminary analysis, we found that five-dimensional RW embeddings offer greater robustness than the ten-dimensional version.
@@ Appendix B In-distribution experimental results discussion
Figure 6 provides a comprehensive visualization of ID test accuracy results across the six benchmark datasets. This analysis complements our OOD findings by validating that high OOD performance does not come at the expense of staying competitive, or even superior, under standard ID conditions. These results also appear in Table 1. On CMNIST, GPS improves the vGIN results significantly across all DG algorithms, while MHA notably decreases ID accuracy. On Motif tasks, backbones overall perform comparably. On SST2, vGIN slightly outperforms GPS, while MHA shows lower performance. On the HIV tasks, vGIN generally achieves the best performance, although GPS is comparable in certain cases. These findings are crucial for practical applications where both ID and OOD samples are encountered, as they show that the superior OOD generalization capabilities of GPS hybrid do not come at the expense of standard ID performance.